# Modeling $(\infty,1)$-categories with Segal spaces

### v1

@@ Abstract.
In this paper, we construct a model structure for $(\infty,1)$-categories on the category of simplicial spaces, whose fibrant objects are the Segal spaces. In particular, we show that it is Quillen equivalent to the models of $(\infty,1)$-categories given by complete Segal spaces and Segal categories. We furthermore prove that this model structure has desirable properties: it is cartesian closed and left proper. As applications, we get a simple description of the inclusion of categories into $(\infty,1)$-categories and of homotopy limits of $(\infty,1)$-categories
@@ 1. Introduction
To address these issues, Rezk developed in [Rez01] an alternative model of $(\infty,1)$-categories using simplicial objects valued in spaces rather than sets. This model is given by the complete Segal spaces and is obtained as a simplicial localization of the model structure on simplicial objects in spaces. Here, a Segal space is a simplicial space $X\colon\Delta^{\mathrm{op}}\to\mathrm{sSet}$ that satisfies the Segal condition, i.e., the Segal map $X_{n}\xrightarrow{\simeq}X_{1}\times_{X_{0}}\cdots\times_{X_{0}}X_{1}$ is a weak homotopy equivalence for all $n\geq 2$. Thinking of $X_{0}$ and $X_{1}$ as spaces of objects and morphisms, the Segal condition allows one to think of $X$ as encoding the algebraic structure of an $(\infty,1)$-category: it identifies each $X_{n}$ as the space of $n$ composable morphisms in $X$, so that the simplicial structure maps of $X$ describe the homotopy coherent composition of morphisms. The additional completeness condition asserts that the space of objects $X_{0}$ is weakly homotopy equivalent to the underlying $\infty$-groupoid of $X$. This condition is somewhat different from the Segal condition: it does not really encode any algebraic structure, but instead ensures that the weak equivalences between complete Segal spaces coincide with homotopical analogues of fully faithful and essentially surjective functors, typically referred to as Dwyer–Kan equivalences. In fact, weak equivalences between Segal spaces are already the Dwyer–Kan equivalences, as any Segal space can be replaced by a Dwyer–Kan equivalent complete Segal space.
@@
Is there a model of $(\infty,1)$-categories in which the fibrant objects are the Segal spaces and the weak equivalences between fibrant objects are the Dwyer–Kan equivalences? [^1] This question seems to have already been asked 14 years ago in a MathSciNet post by Schommer-Pries: https://mathoverflow.net/questions/29728/a-model-category-of-segal-spaces
A first attempt at answering this question was given by replacing the completeness condition with a discreteness condition, requiring the space of objects to be a set. This was formalized through a model structure, first constructed by Pellissier [Pel02] and further studied by Bergner [Ber07a, Ber07b] and Simpson [Sim12], in which the fibrant objects are the Segal categories, i.e., the simplicial spaces $X\colon\Delta^{\mathrm{op}}\to\mathrm{sSet}$ that satisfy the Segal condition and such that the space of objects $X_{0}$ is a set. Notably, since the requirement that $X_{0}$ be a set is not a homotopical condition, the model structure cannot be defined on the whole category of simplicial spaces. Instead, it must be restricted to the full subcategory $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ consisting of those simplicial spaces whose level 00 is a set, referred to as precategory objects in $\mathrm{sSet}$. However, this discreteness condition is not satisfied by all of the above examples, e.g. not by the $(\infty,1)$-category of bordisms, and it is not always convenient to choose a set of vertices in the space of objects of a Segal space.
In this paper, we provide a positive answer to the above question. We construct a model structure on the category of simplicial spaces, where the fibrant objects are the Segal spaces, and the “completeness” or “discreteness” condition is transferred to the cofibrant objects. Specifically, the cofibrant objects are defined to be the simplicial spaces $X\colon\Delta^{\mathrm{op}}\to\mathrm{sSet}$ such that the space of objects $X_{0}$ is weakly homotopy equivalent to a set. The following result is proven in Section 3 and is stated as Theorem 3.1.
@@
This model structure is very similar to the categorical model structure on categories: the cofibrations, fibrations, and weak equivalences between Segal spaces are homotopical analogues of functors that are injective on objects, isofibrations, and equivalences of categories, respectively. Furthermore, it provides, as desired, a model of $(\infty,1)$-categories, as it is Quillen equivalent to the complete Segal space model structure. In fact, it sits somewhere between Segal categories and complete Segal spaces, by the following combination of Theorems 4.4, 4.9 and 4.13.
@@
Moreover, the categorical model structure retains the desirable properties of the model structure for complete Segal spaces. The following result is a combination of Theorems 5.1 and 5.4.
@@
These properties of the categorical model structure often allow one to deal with Segal spaces directly, without having to complete them. As a first example, recall that the nerve of a category defines a (discrete) Segal space which is generally not complete. As a consequence, the nerve does not provide an appropriate homotopical functor into the complete Segal space model structure. However, the nerve does induce a right Quillen functor from the categorical model structure on categories to the model structure on simplicial spaces from A, as we show in Proposition 6.1. As a second example, the categorical model structure can be used to compute pullbacks of $(\infty,1)$-categories—and more generally any limit of $(\infty,1)$-categories—modeled by non-complete Segal spaces, in a way similar to the computation of homotopy pullbacks of ordinary categories. Explicitly, we prove in Proposition 6.2 that the pullback of a map between Segal spaces along an isofibration is already a homotopy pullback in the complete Segal space model structure.
@@ Acknowledgments
During the realization of this work, the first author was a member of the Collaborative Research Centre “SFB 1085: Higher Invariants” funded by the DFG. The second author was supported by the CNRS, under the program Projet Exploratoire de Premier Soutien “Jeune chercheuse, jeune chercheur” (PEPS JCJC).
@@ 2. Segal spaces, Dwyer–Kan equivalences, and isofibrations
In this section, we provide the necessary background on Segal spaces and $(\infty,1)$-categorical notions of equivalences and fibrations between them. In Section 2.1, we recall the notion of Segal spaces, as well as their associated model structure on simplicial spaces. Next, in Section 2.2, we recall the definition of Dwyer–Kan equivalences between Segal spaces and establish a useful lemma about them. Finally, in Section 2.3, we introduce a notion of isofibrations between Segal spaces and prove some technical results related to them.
@@
For $m\geq 0$, we will write $F[m]\colon\Delta^{\mathrm{op}}\to\mathrm{sSet}$ for the representable in the categorical direction, sending an object $[k]\in\Delta$ to the constant simplicial set at the set $\Delta([k],[m])$. We denote by $\partial F[m]$ its boundary and by $\mathrm{Sp}[m]$ its spine.
@@ 2.2. Dwyer–Kan equivalences
[…]
@@
[…]
@@
[…]
@@
[…]
The horizontal maps are fibrations in $\mathrm{sSet}$ since $X$ and $Y$ are Reedy fibrant. Consequently, condition (iii) is equivalent to the total rectangle being a homotopy pullback and condition (iv) is equivalent to the left square being a homotopy pullback. We will prove this by induction on $m$, the initial case $m=1$ being precisely (ii). The composite of the two leftmost squares is a homotopy pullback, since the top and bottom maps are both weak equivalences in $\mathrm{sSet}$ by the Segal conditions on $X$ and $Y$. Furthermore, the right-hand square and the composition of the two rightmost squares are both homotopy pullback squares. Indeed, this follows from the fact that the monomorphisms
@@
Every weak equivalence $f\colon X\to Y$ between Segal spaces in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ is a Dwyer–Kan equivalence. Indeed, we can write $f$ as the composition of a Reedy trivial fibration $p\colon X^{\prime}\to Y$ and a section of a Reedy trivial fibration $i\colon X\to X^{\prime}$. Since $p$ is surjective on objects, it is a Dwyer–Kan equivalence by Lemma 2.11. The map $i$ is a Dwyer–Kan equivalence by 2-out-of-3.
@@ 2.3. Isofibrations
Finally, let us introduce the following version of isofibrations between Segal spaces. For this we first recall that classically, an isofibration is a functor with the right lifting property against either inclusion $[0]\hookrightarrow I[1]$, where $I[1]$ denotes the free-living isomorphism.
@@
Assertions (i) and (ii) follow from the fact that each inner horn inclusion $L^{k}[m]\hookrightarrow F[m]$ in the categorical direction is a trivial cofibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ by [JT07, Lemma 3.5].
For (iii), using Lemma 2.10, it suffices to show that the inclusion $X_{-,0}\to RX$ induces an isomorphism on categorifications. For this, recall from [Joy08, Proposition 1.11] that the category $c(X_{-,0})$ has set of objects $X_{0,0}$, and hom set at $x,y\in X_{0,0}$ given by the quotient of $\mathrm{map}_{X}(x,y)_{0}$ by the following equivalence relation: $\alpha\sim\beta$ if there exists $h\in X_{2,0}$ such that $d_{2}h=\alpha$, $d_{1}h=\beta$, and $d_{0}h=s_{0}y$. Unraveling the definitions, the functor $c(X_{-,0})\to c(RX)$ is now given by the identity on objects and, on hom sets, by the canonical map
@@
$\partial F[2]\times\Delta[1]\amalg_{\partial F[2]\times\{0\}}F[2]\times\{0\}$$X$$F[2]\times\Delta[1]$$H$$x$$y$$y$$x$$y$$y$$\alpha$$\alpha$$\beta$$\alpha$$h^{\prime}$
@@
The map $f\colon X\to Y$ has the right lifting property against $F[0]\hookrightarrow NI[1]$ if and only if the induced map $f_{-,0}\colon X_{-,0}\to Y_{-,0}$ has the right lifting property against $F[0]\hookrightarrow NI[1]$. The map $f_{-,0}$ is an inner fibration between quasi-categories by Lemma 2.14 (ii), so that [Joy02, Proposition 2.4] implies that the map $f_{-,0}$ has the said lifting property if and only if the induced functor $c(X_{-,0})\to c(Y_{-,0})$ between homotopy categories is an isofibration. By Lemma 2.14 (iii), we conclude that this is the case if and only if $f$ is an isofibration. ∎
Finally, we prove a technical result about isofibrations, that will be useful later on.
@@
To see that the outer and right-hand square are pullbacks, note that the pullback of the given cospan is the space of maps $(\partial)F[m]\to X$ whose restriction to $\mathrm{Sp}[m]$ classifies $m$ composable morphisms in $X$ that all induce isomorphisms in $\mathrm{ho}X$. Any composition of these morphisms defines an isomorphism in $\mathrm{ho}X$ as well, so that any composite $F[1]\to(\partial)F[m]\to X$ defines an element in $X_{\mathrm{hoeq}}$. This shows that the pullback is isomorphic to $(X^{\simeq})_{m}$ or $\mathrm{Map}(\partial F[m],X^{\simeq})$. The left-hand square is then also a pullback square by cancellation. Using this, the fact that $X$ is a Segal space implies that $X^{\simeq}$ is also a Segal space. By Lemma 2.14 (i), we get that $(X^{\simeq})_{-,0}$ is a quasi-category, whose homotopy category coincides with $\mathrm{ho}(X^{\simeq})\subseteq\mathrm{ho}X$ by Lemma 2.14 (iii). Every morphism in the quasi-category $(X^{\simeq})_{-,0}$ is therefore invertible, so that $(X^{\simeq})_{-,0}$ is a Kan complex by [Joy02, Corollary 1.4].
For (iv), using (iii), it suffices to show that $(X^{\simeq})_{-,0}\to(Y^{\simeq})_{-,0}$ is an inner fibration which lifts against $F[0]\hookrightarrow NI[1]$. Using the above left pullback square, one can show that the $m$-th relative matching map of $X^{\simeq}\to Y^{\simeq}$ is a pullback of the $m$-th relative matching map of $X\to Y$. The fact that $X\to Y$ is a Reedy fibration then implies that ${X^{\simeq}\to Y^{\simeq}}$ is also a Reedy fibration. By Lemma 2.14 (ii), we get that $(X^{\simeq})_{-,0}\to(Y^{\simeq})_{-,0}$ is an inner fibration. Using [Joy02, Proposition 2.4] and the isomorphisms $\mathrm{ho}((X^{\simeq})_{-,0})\cong\mathrm{ho}(X^{\simeq})$ and $\mathrm{ho}((Y^{\simeq})_{-,0})\cong\mathrm{ho}(Y^{\simeq})$ from Lemma 2.14 (iii), it suffices to show that $X^{\simeq}\to Y^{\simeq}$ has the right lifting property against $F[0]\hookrightarrow NI[1]$. By Proposition 2.15 and (i), this follows from the fact that $X\to Y$ is an isofibration. ∎
@@
Using Lemma 2.17 (i) and (ii), this is equivalent to showing that $X^{\simeq}\to Y^{\simeq}$ has the right lifting property against the above map. This follows from the fact that $(X^{\simeq})_{-,0}\to(Y^{\simeq})_{-,0}$ is fibration in $\mathrm{sSet}$ by Lemma 2.17 (iv).

### v2

@@ Abstract.
In this paper, we construct a model structure for $(\infty,1)$-categories on the category of simplicial spaces, whose fibrant objects are the Segal spaces. In particular, we show that it is Quillen equivalent to the models of $(\infty,1)$-categories given by complete Segal spaces and Segal categories. We furthermore prove that this model structure has desirable properties: it is cartesian closed and left proper. As applications, we get a simple description of the inclusion of categories into $(\infty,1)$-categories and of homotopy limits of $(\infty,1)$-categories.
1991 Mathematics Subject Classification:
18N60, 55U35, 18N50
@@ 1. Introduction
An alternative model of $(\infty,1)$-categories, using simplicial objects valued in spaces rather than sets, has been developed by Rezk [Rez01]. This model is given by the complete Segal spaces and is obtained as a simplicial localization of the model structure on simplicial objects in spaces. Here, a Segal space is a simplicial space $X\colon\Delta^{\mathrm{op}}\to\mathrm{sSet}$ that satisfies the Segal condition, i.e., the Segal map $X_{n}\xrightarrow{\simeq}X_{1}\times_{X_{0}}\cdots\times_{X_{0}}X_{1}$ is a weak homotopy equivalence for all $n\geq 2$. Thinking of $X_{0}$ and $X_{1}$ as spaces of objects and morphisms, the Segal condition allows one to think of $X$ as encoding the algebraic structure of an $(\infty,1)$-category: it identifies each $X_{n}$ as the space of $n$ composable morphisms in $X$, so that the simplicial structure maps of $X$ describe the homotopy coherent composition of morphisms. The additional completeness condition asserts that the space of objects $X_{0}$ is weakly homotopy equivalent to the underlying $\infty$-groupoid of $X$. This condition is somewhat different from the Segal condition: it does not really encode any algebraic structure, but instead ensures that the weak equivalences between complete Segal spaces coincide with homotopical analogues of fully faithful and essentially surjective functors, typically referred to as Dwyer–Kan equivalences. In fact, weak equivalences between Segal spaces are already the Dwyer–Kan equivalences, as any Segal space can be replaced by a Dwyer–Kan equivalent complete Segal space.
@@
Is there a model of $(\infty,1)$-categories in which the fibrant objects are the Segal spaces and the weak equivalences between fibrant objects are the Dwyer–Kan equivalences? [^1] This question was asked 14 years ago in a MathOverflow post by Chris Schommer-Pries: https://mathoverflow.net/questions/29728/a-model-category-of-segal-spaces
One possible way to address this question is by restricting to a subcategory of simplicial spaces and modeling $(\infty,1)$-categories with Segal categories. These were originally introduced by Hirschowitz and Simpson [HS98] and in fact predate the notion of complete Segal spaces. They arise as the fibrant objects of a model structure, first constructed by Pellissier [Pel02] and further studied by Bergner [Ber07a, Ber07b] and Simpson [Sim12]. In this model, the completeness condition on a Segal space is replaced by a discreteness condition, requiring the space of objects to be a set. Since this is notably not a homotopical condition, this model structure can only be defined on the full subcategory $\mathcal{P}\mathrm{Cat}(\mathrm{sSet})$ consisting of those simplicial spaces whose level 0 is a set, referred to as precategory objects in $\mathrm{sSet}$. However, this discreteness condition is not satisfied by all of the above examples, e.g. not by the $(\infty,1)$-category of bordisms, and it is not always convenient to choose a set of vertices in the space of objects of a Segal space.
In this paper, we provide a positive answer to the above question. We construct a model structure on the category of simplicial spaces, where the fibrant objects are the Segal spaces, and the “completeness” or “discreteness” condition is transferred to the cofibrant objects. Specifically, the cofibrant objects are defined to be the simplicial spaces $X\colon\Delta^{\mathrm{op}}\to\mathrm{sSet}$ such that the space of objects $X_{0}$ is weakly homotopy equivalent to a set. The following result is proven in Section˜3 and is stated as Theorem˜3.1.
@@
This model structure is very similar to the categorical model structure on categories: the cofibrations, fibrations, and weak equivalences between Segal spaces are homotopical analogues of functors that are injective on objects, isofibrations, and equivalences of categories, respectively. Furthermore, it provides, as desired, a model of $(\infty,1)$-categories, as it is Quillen equivalent to the complete Segal space model structure. In fact, it sits somewhere between Segal categories and complete Segal spaces, by the following combination of Theorems˜4.4, 4.9 and 4.13.
@@
Moreover, the categorical model structure retains the desirable properties of the model structure for complete Segal spaces. The following result is a combination of Theorems˜5.1 and 5.4.
@@
These properties of the categorical model structure often allow one to deal with Segal spaces directly, without having to complete them. As a first example, recall that the nerve of a category defines a (discrete) Segal space which is generally not complete. As a consequence, the nerve does not provide an appropriate homotopical functor into the complete Segal space model structure. However, the nerve does induce a right Quillen functor from the categorical model structure on categories to the model structure on simplicial spaces from ˜A, as we show in Proposition˜6.1. As a second example, the categorical model structure can be used to compute pullbacks of $(\infty,1)$-categories—and more generally any limit of $(\infty,1)$-categories—modeled by non-complete Segal spaces, in a way similar to the computation of homotopy pullbacks of ordinary categories. Explicitly, we prove in Proposition˜6.2 that the pullback of a map between Segal spaces along an isofibration is already a homotopy pullback in the complete Segal space model structure.
@@ Acknowledgments
We are grateful to João Candeias and to the referee for their careful reading of the paper and their useful comments. During the realization of this work, the first author was a member of the Collaborative Research Centre “SFB 1085: Higher Invariants” funded by the DFG. The second author was supported by the CNRS, under the program Projet Exploratoire de Premier Soutien “Jeune chercheuse, jeune chercheur” (PEPS JCJC).
@@ 2. Segal spaces, Dwyer–Kan equivalences, and isofibrations
In this section, we provide the necessary background on Segal spaces and $(\infty,1)$-categorical notions of equivalences and fibrations between them. In Section˜2.1, we recall the notion of Segal spaces, as well as their associated model structure on simplicial spaces. Next, in Section˜2.2, we recall the definition of Dwyer–Kan equivalences between Segal spaces and establish a useful lemma about them. Finally, in Section˜2.3, we introduce a notion of isofibrations between Segal spaces and prove some technical results related to them.
@@
For $m\geq 0$, we will write $F[m]\colon\Delta^{\mathrm{op}}\to\mathrm{sSet}$ for the representable in the categorical direction, sending an object $k]\in\Delta$ to the constant simplicial set at the set $\Delta([k],[m])$. We denote by $\partial F[m]$ its boundary and by $\mathrm{Sp}[m]$ its spine.
@@ 2.2. Dwyer–Kan equivalences
$c\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}\rightleftarrows\mathrm{Cat}\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0muN$
@@
We write $(-)_{0}\colon\mathrm{sSet}^{\Delta^{\mathrm{op}}}\rightleftarrows\mathrm{sSet}\penalty 10000\mskip 6.0mu plus 1.0mu\mathpunct{}\nonscript\mkern-3.0mu{:}\mskip 2.0mu\mathrm{cosk}_{0}$ for the adjunction whose left adjoint sends a simplicial space $X$ to its underlying space $X_{0}$, and whose right adjoint sends $K\in\mathrm{sSet}$ to the simplicial space $\mathrm{cosk}_{0}(K)$ given at $m\geq 0$ by $\mathrm{cosk}_{0}(K)_{m}\cong K^{\times(m+1)}$.
@@
$X_{1}$$Y_{1}$$X_{0}\times X_{0}$$Y_{0}\times Y_{0}$$\langle 0,1\rangle^{*}$$\langle 0,1\rangle^{*}$
@@
$X_{m}$$\mathrm{Map}(\partial F[m],X)$$\mathrm{Map}(\mathrm{Sp}[m],X)$$X_{0}^{\times(m+1)}$$Y_{m}$$\mathrm{Map}(\partial F[m],Y)$$\mathrm{Map}(\mathrm{Sp}[m],Y)$$Y_{0}^{\times(m+1)}$
The horizontal maps are fibrations in $\mathrm{sSet}$ since $X$ and $Y$ are Reedy fibrant. Consequently, condition (iii) is equivalent to the total rectangle being a homotopy pullback and condition (iv) is equivalent to the left square being a homotopy pullback. We will prove these are homotopy pullbacks by induction on $m$, the initial case $m=1$ being precisely (ii). The composite of the two leftmost squares is a homotopy pullback, since the top and bottom maps are both weak equivalences in $\mathrm{sSet}$ by the Segal conditions on $X$ and $Y$. Furthermore, the right-hand square and the composition of the two rightmost squares are both homotopy pullback squares. Indeed, this follows from the fact that the monomorphisms
@@
Every weak equivalence $f\colon X\to Y$ between Segal spaces in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ is a Dwyer–Kan equivalence. Indeed, we can write $f$ as the composition of a Reedy trivial fibration $p\colon X^{\prime}\to Y$ and a section of a Reedy trivial fibration $i\colon X\to X^{\prime}$. Since $p$ is surjective on objects, it is a Dwyer–Kan equivalence by Lemma˜2.11. The map $i$ is a Dwyer–Kan equivalence by 2-out-of-3.
@@ 2.3. Isofibrations
Finally, let us introduce the following version of isofibrations between Segal spaces. For this we first recall that classically, an isofibration is a functor with the right lifting property against either inclusion $0]\hookrightarrow I[1]$, where $I[1]$ denotes the free-living isomorphism.
@@
Assertions (i) and (ii) follow from the fact that each inner horn inclusion $L^{k}[m]\hookrightarrow F[m]$ for $0<k<m$, where $L^{k}[m]$ denotes the $k$-th horn, in the categorical direction is a trivial cofibration in $\mathrm{sSet}^{\Delta^{\mathrm{op}}}_{\mathrm{Seg}}$ by [JT07, Lemma 3.5].
For (iii), using Lemma˜2.10, it suffices to show that the inclusion $X_{-,0}\to RX$ induces an isomorphism on categorifications. For this, recall from [Joy08, Proposition 1.11] that the category $c(X_{-,0})$ has set of objects $X_{0,0}$, and hom set at $x,y\in X_{0,0}$ given by the quotient of $\mathrm{map}_{X}(x,y)_{0}$ by the following equivalence relation: $\alpha\sim\beta$ if there exists $h\in X_{2,0}$ such that $d_{2}h=\alpha$, $d_{1}h=\beta$, and $d_{0}h=s_{0}y$. Unraveling the definitions, the functor $c(X_{-,0})\to c(RX)$ is now given by the identity on objects and, on hom sets, by the canonical map
@@
[…]
@@
The map $f\colon X\to Y$ has the right lifting property against $F[0]\hookrightarrow NI[1]$ if and only if the induced map $f_{-,0}\colon X_{-,0}\to Y_{-,0}$ has the right lifting property against $F[0]\hookrightarrow NI[1]$ (viewed as a map of simplicial sets). The map $f_{-,0}$ is an inner fibration between quasi-categories by Lemma˜2.14 (ii), so that [Joy02, Proposition 2.4] implies that the map $f_{-,0}$ has the said lifting property if and only if the induced functor $c(X_{-,0})\to c(Y_{-,0})$ between homotopy categories is an isofibration. By Lemma˜2.14 (iii), we conclude that this is the case if and only if $f$ is an isofibration. ∎
Finally, we prove a technical result about isofibrations that will be useful later on.
@@
To see that the outer and right-hand square are pullbacks, note that the pullback of the given cospan is the space of maps $(\partial)F[m]\to X$ whose restriction to $\mathrm{Sp}[m]$ classifies $m$ composable morphisms in $X$ that all induce isomorphisms in $\mathrm{ho}X$. Any composition of these morphisms defines an isomorphism in $\mathrm{ho}X$ as well, so that any composite $F[1]\to(\partial)F[m]\to X$ defines an element in $X_{\mathrm{hoeq}}$. This shows that the pullback is isomorphic to $(X^{\simeq})_{m}$ or $\mathrm{Map}(\partial F[m],X^{\simeq})$. The left-hand square is then also a pullback square by cancellation. Using this, the fact that $X$ is a Segal space implies that $X^{\simeq}$ is also a Segal space. By Lemma˜2.14 (i), we get that $(X^{\simeq})_{-,0}$ is a quasi-category, whose homotopy category coincides with $\mathrm{ho}(X^{\simeq})\subseteq\mathrm{ho}X$ by Lemma˜2.14 (iii). Every morphism in the quasi-category $(X^{\simeq})_{-,0}$ is therefore invertible, so that $(X^{\simeq})_{-,0}$ is a Kan complex by [Joy02, Corollary 1.4].
For (iv), using (iii), it suffices to show that $(X^{\simeq})_{-,0}\to(Y^{\simeq})_{-,0}$ is an inner fibration which lifts against $F[0]\hookrightarrow NI[1]$. Using the above left pullback square, one can show that the $m$-th relative matching map of $X^{\simeq}\to Y^{\simeq}$ is a pullback of the $m$-th relative matching map of $X\to Y$. The fact that $X\to Y$ is a Reedy fibration then implies that $X^{\simeq}\to Y^{\simeq}$} is also a Reedy fibration. By Lemma˜2.14 (ii), we get that $(X^{\simeq})_{-,0}\to(Y^{\simeq})_{-,0}$ is an inner fibration. Using [Joy02, Proposition 2.4] and the isomorphisms $\mathrm{ho}((X^{\simeq})_{-,0})\cong\mathrm{ho}(X^{\simeq})$ and $\mathrm{ho}((Y^{\simeq})_{-,0})\cong\mathrm{ho}(Y^{\simeq})$ from Lemma˜2.14 (iii), it suffices to show that $X^{\simeq}\to Y^{\simeq}$ has the right lifting property against $F[0]\hookrightarrow NI[1]$. By Proposition˜2.15 and (i), this follows from the fact that $X\to Y$ is an isofibration. ∎
@@
Using Lemma˜2.17 (i) and (ii), this is equivalent to showing that $X^{\simeq}\to Y^{\simeq}$ has the right lifting property against the above map. This follows from the fact that $(X^{\simeq})_{-,0}\to(Y^{\simeq})_{-,0}$ is fibration in $\mathrm{sSet}$ by Lemma˜2.17 (iv).